Fruit drink a consists of 6% fruit juice and drink. B contains 15% fruit juice and drink. how much of each kind of drink should you mix to get 4L of 10% fruit juice.

i tried x+y=4 and .06x+.15y=.4 for my system but i cant get the right answers when i check them

Answer by lwsshak3(11628)

Fruit drink a consists of 6% fruit juice and drink. B contains 15% fruit juice and drink. how much of each kind of drink should you mix to get 4L of 10% fruit juice.
..
let x=liters of 6% drink to be mixed
4-x=liters of 15% drink to be mixed
.06x+.15(4-x)=.10(4)
.06x+.60-.15x=.40
.09x=.20
x=.20/.09=2.22 liters
4-x=1.78 liters
ans:
2.22 liters of 6% drink should be mixed with 1.78 liters of 15% drink to get 4 liters of 10% drink.
Looks like you had the correct equations for a solution.